OpenStudy (jade27p):

Describe how to transform the quantity of the third root of x to the fourth power, to the fifth powerinto an expression with a rational exponent. Make sure you respond with complete sentences. @phi

OpenStudy (jade27p):

can you check my work

OpenStudy (jade27p):

x/a^n=x 20/3

OpenStudy (jade27p):

do i simplify the radical next

OpenStudy (phi):

if you mean \[ \left( \sqrt[3]{x^4}\right)^5 = x^\frac{20}{3} \] yes, that is correct. I don't know what a^n means in x/a^n=x 20/3

OpenStudy (jade27p):

idk how to type the square root on my laptop

OpenStudy (jade27p):

but yes thats what i have on my paper

OpenStudy (phi):

the "cube root" of x^4 can be written as \[ \left(x^4\right)^\frac{1}{3} \] and we can re-write that by "multiplying the exponents" to get \[ x^\frac{4}{3} \] and if we raise that to the 5th power \[ \left(x^\frac{4}{3} \right)^5 \] we multiply by expoents again to simplify to the answer

OpenStudy (jade27p):

is that x^3

OpenStudy (phi):

are you asking about \[ \left(x^\frac{4}{3} \right)^5 \]? to simplify to just one expoent , figure out 5* 4/3

OpenStudy (jade27p):

20/3

OpenStudy (phi):

yes, so x^(20/3) is another way to write \[ \left( \sqrt[3]{x^4}\right)^5 \]

OpenStudy (jade27p):

so these are the steps

OpenStudy (phi):

yes, the first step is to replace the cube root "sign" with an exponent of 1/3 see the post a few steps up.